OpenStudy (anonymous):

Let f(x) = x - 3 and g(x) = x + 11. Find f(x) * g(x). Answer Options: x^2 - 8x - 33 x^2 - 8x +14 x^2 + 8x +14 x^2 + 8x -33

OpenStudy (anonymous):

=(x-3)(x+11) =x^2 +11 x -3 x + 10 =x^2 + 8 x + 10

OpenStudy (anonymous):

Sorry x^2 + 8 x + 10

OpenStudy (anonymous):

That's not in the answer choices though

OpenStudy (anonymous):

Sorry =x2+11x-3x - 33 =x^2 + 8x -33

OpenStudy (anonymous):

That's it girl

OpenStudy (anonymous):

Haha it's okay , I'm understanding a little but though . So thank you

OpenStudy (anonymous):

Could you help me with one more ?

OpenStudy (anonymous):

Let me do it over it again for you (x-3)(x+11) Multiply the first two in bracket x^2+11x Then multiply the second two -3x - 33 so its x^2+ 11x - 3x -33 x^2 + 8x -33

OpenStudy (anonymous):

Thanks !

OpenStudy (anonymous):

ya welcome